Translate to a quadratic equation, then solve using the quadratic formula.
The length of a rectangular air filter is 3 inches less than twice the width. Find the length and the width of the filter if the area is 665 square inches.
What are the length and the width??

Answer by mananth(16946)

width = w
length = 2w-3
Area =l*w
w(2w-3)=665
2w^2-3w-665=0
2w62-38W+35w-665=0
2w(w-19)+35(w-19)=0
(w-19)(2w+35)=0
w=19 OR -35/2
ignore negative

width = 19 inches
length = 2*19-3
=35 inches